### v4

# High-throughput Parasitic-independent Probe Thermal Resistance Calibration for Robust Thermal Mapping with Scanning Thermal Microscopy

Abstract
Nanostructured materials, critical for thermal management in semiconductor devices, exhibit a strong size dependence in thermal transport. Studying thermal resistance variation across grain boundaries is critical for designing effective thermal interface materials. Frequency-domain Thermoreflectance (FDTR)-based techniques can provide thermal resistance mapping at the micrometer (m) scale. Scanning Thermal Microscopy (SThM) enables quantification of local thermal transport with significantly higher spatial resolution ( nm). However, challenges in quantifying the raw signal to thermal conductivity and surface sensitivity limit its widespread adoption for understanding nanoscale heat transport and defect-mediated thermal properties in nanostructured films. Here, we introduce a circuit-based probe thermal resistance () calibration technique independent of parasitic heat pathways, enabling accurate determination of probe heat dissipation and tip temperature rise, thereby allowing extraction of local thermal resistance. SThM achieved sub-100 nm spatial resolution in mapping thermal resistance across a 15 nm-thick Al film deposited via e-beam evaporation on SiO2 substrate. The thermal resistance maps are converted to thermal conductivity using robust analytical and finite element models that account for tip-sample geometry, lateral heat spreading, and buried interface effects. Gaussian distribution fitting of pixel-level thermal resistance values yields W/(mK) for the ultra-thin Al film (13–15 nm), representing a 5.3-fold reduction from bulk aluminum (237 W/(mK)). These results agree with published experimental data and theoretical frameworks explaining thickness-dependent heat transport in ultra-thin metallic films.
I Introduction
Across the last two decades, continuous materials scaling coupled with precise nanostructuring has powered a remarkable run of progress in electronic10019538; munde20253d; banerjee20023, photonicostroverkhova2016organic; huang2020recent, and quantum technologiesputranto2024deep, reinforced by advances in MEMS/NEMS zhu2019development; lyshevski2018mems, nanoscale sensors for biomedicinehannah2020developments; wang2013advances; tovar2023recent, and scanning-probe/optical instrumentationcahill2003nanoscale. Miniaturization fundamentally modifies transport, mechanical, and interfacial behavior. As a result, size-dependent variations in material properties become critical considerations in modeling, metrology, and device design. For example, silicon (Si) thermal conductivity () is reduced from its bulk value of 138 W/mK to 2.2 W/mK for a 155 nm Si thin filmmcconnell2005thermal; huang2019thermal. When the physical dimension becomes comparable to or smaller than the mean free path () of heat carriers (electrons and phonons), the macroscopic equilibrium and continuum assumptions break down. Therefore, understanding the nanoscale energy transport phenomena, both analytically and experimentally, is crucial for the continued miniaturization of technology.
Thermal metrology plays a vital role in validating theoretical models and advancing the understanding of energy transport mechanisms. Conventional techniques such as -thermocouples, microbridge and suspended devices (), infrared thermography (IR), and liquid crystal thermography are used to measure macroscopic thermal properties of devices. Recently, with technological advancements in device fabrication and deposition processes, the technique has been extended for use with thin films and microscale devices by fabricating bridges a few thick cahill1987thermal. This method enables accurate characterization of both in-plane and cross-plane thermal transport in metallic, semiconducting, and dielectric thin films. A major limitation of the 3 technique is its requirement for complex sample preparation and the complexities arising from voltage oscillations when measuring metals and semiconductors, thus limiting its applicability to a narrow range of materialscahill2002thermometry. Optical based techniques such as frequency- and time-domain thermoreflectance spectroscopy (FDTR and TDTR), Raman spectroscopy, laser flash techniques, and fluorescence techniques are widely used for measuring the temperature and thermal transport properties of microstructured materials christofferson2008microscale; tiwari2013anisotropic; zeng2017photothermal. In contrast to the method, optical-based techniques such as TDTR and FDTR enable non-contact characterization of thermal properties, minimizing physical interference with the sample. Recent advancements in laser source technology have significantly enhanced the capability of TDTR and FDTR techniques to characterize the thermal properties of materials with superior temporal and spatial resolution isotta2023microscale. Depending on the type of laser used in the measurement, these methods can achieve temporal resolutions ranging from a few nanoseconds to 30 fs (1 fs = sec). It is worth noting that the electronic relaxation time () typically lies in the range of about 10–50 fs. Therefore, TDTR and FDTR are useful techniques for measuring the thermal transport properties associated with electron-phonon coupling mechanisms. However, these optical-based thermal measurement methods suffer from an intrinsic spatial resolution limitation imposed by the pump–probe spot size (typically 0.5–10 m), rendering them inadequate for reliably imaging thermal heterogeneities arising from nanoscale defects.
In the late 1990s, the development of scanning probe microscopy (SPM)–based techniques, such as scanning thermal microscopy (SThM) and near-field scanning optical microscopy (NSOM) provided unprecedented spatial resolution for nanoscale characterization and manipulation. In general, NSOM requires optically compatible and flat samples, is limited by weak signal-to-noise ratio, and exhibits greater tip fragility compared to SThM. Table 1 summarizes the quantitative capabilities of thermal metrology techniques for microelectronics. Current commercially available SThMs are based on the working principle designed by Majumdar et al.majumdar1993thermal as illustrated in Fig. 1(a). The SThM system uses specialized thermal probes (TPs) with tip radii below 100 nm, capable of resolving temperature differences smaller than 0.1 with sub-100 nm spatial resolution, thereby enabling quantitative nanoscale thermal characterization of materials. SThM can operate in either temperature contrast mode (TCM) or conductivity contrast mode (CCM). In TCM, a tip-integrated thermocouple measures the junction temperature at the tip-sample interface, whereas the CCM mode provides qualitative mapping of local thermal conductivity across a broad range of materials, including metallic, semiconducting, and dielectric thin films. Transforming these qualitative signals into quantitative thermal conductivity values is not trivial because of the complex coupled thermal transport at the tip-sample interface. There is no standardized technique for obtaining quantitative measurements of thermal resistance with SThM. Recently, various theoretical modeling, calibration strategies, and instrument add-on techniques have been reportedgucmann2021scanning; li2022comprehensive; li2025quantitative. Bodzenta et al. and Li et al. provided comprehensive reviews of SThM theoretical models and experimental strategies for quantitative thermal measurementsbodzenta2022scanning; li2025quantitative.
In this work, we present a circuit-based calibration technique for measuring non-contact thermal resistance (), also known as probe thermal resistance. Compared with conventional thermal-stage methods, this calibration approach reduces sensitivity to temperature-controller errors and environmental factors (e.g., moisture), and is less affected by common artifacts, including calibration-sample variations. First, we construct thermal-resistance maps across the film surface by coupling scanning thermal measurements with finite-element heat-transfer modeling of the tip–sample mechanical contact. These maps resolve spatial variations in thermal transport that arise from differences in crystallite size and orientation, effects that are most pronounced in polycrystalline and amorphous films. Because the maps carry the imprint of surface preparation and condition—the same steps a film undergoes in routine device-integration flows—the resulting resistance estimates reflect the as-processed device film. Building on the same calibrated contact model, we then extract the film’s effective thermal conductivity () using an experimentally anchored procedure that combines accurate instrument calibration with contact heat-transfer modeling. Sec. II of this paper discusses the theoretical modeling of SThM required for understanding different heat transfer mechanisms at the tip-sample contact. Sec. III describes the methodology used to extract of samples and the post-processing steps to study the underlying nanoscale thermal transport physics and to map the corresponding thermal resistances. In Sec. IV, we present the results obtained through the techniques described in Sec.s II and III. Finally, we summarize our results with concluding remarks.
| Technique | Mechanism | Spatial Resolution | Temperature Resolution |
| Infrared thermography | Thermal radiation (Planck, Stefan–Boltzmann) | 3–10 m | 0.02–1 K |
| Photothermal Raman | Raman shift or Stokes/anti-Stokes ratio | 0.5–1 m | 1–10 K |
| 3 method | AC Joule heating with electrical sensing | — | 0.1 mK–1 K |
| TDTR / FDTR | Reflectivity-based transient thermometry | 1–10 m | 0.01 K |
| NSOM | Near-field optical heat detection | 50–100 nm | 0.1–1 K |
| SThM | AFM-integrated thermocouple/Pd thermistor tip | 10–50 nm | 0.1 K |
Data adapted from Refs. christofferson2008microscale; cahill1987thermal; zhang2007nano.
II Background of The Quantitative SThM Measurement Technique
A schematic of the SThM probe with an integrated Wheatstone bridge is shown in Fig. 1(a). The Wheatstone bridge helps detect changes in the thermal resistance of the temperature-sensitive tip material. SThM provides qualitative maps of both the thermal signal and the surface topography of the scanned region, as shown in Fig. 1(b,c). The topography map reveals the nanoscale geometrical features of the surface, whereas the thermal signal maps offer qualitative insights into the local thermal conductivity of the sample with nanoscale spatial resolution. Furthermore, SThM exhibits a high signal-to-noise ratio, enabling more accurate and reliable measurements (see Fig. 1 (d,e)). However, correlating these measured signals with the intrinsic thermal properties of the materials requires a thorough theoretical understanding of the underlying physical mechanisms.
Theoretical modeling links the probe voltage signals measured by the Wheatstone bridge detector to the sample’s thermal properties. Therefore, accurate quantitative measurements of thermal conductivity heavily rely on mathematical modeling to transform probe voltage signals into thermal conductivity values. Thermal transport between the probe and sample involves multiple heat transfer channels, including conduction through the solid–solid contact (tip–sample) and the water meniscus (from the moisture), convection through air, and thermal radiation, which becomes significant at nanometer scale tip–sample distances (see Fig. 2). Each of these channels has its corresponding contribution, which depends on the instrumentation, measurement environment, and sample under consideration. Various analytical and numerical approaches have been reported for studying each heat transfer channel and its significance in quantitative measurements. A recent review by Li et al. comprehensively discusses multiple analytical and numerical approaches for calculating heat transfer through these channels li2025quantitative.
This work uses a thermistor probe tip that consists of Palladium (Pd) resistor material with a thermal coefficient of resistance (TCR) of 3800 , as specified by the manufacturer. Importantly, the TCR value is not affected by the experimental procedure or daily use of the probepernot2021frequency; huang2024violation. Since Pd is a metal, within a relatively small temperature range, the electrical resistance of the tip () varies linearly with temperature (T).
where is the electrical resistance of the tip at temperature and is the TCR. The tip temperature is measured by applying an electric current to the probe and monitoring the change in electrical resistance of the tip with SThM software.
When the probe is in air without contacting the sample, heat is dissipated by conduction through the cantilever, convection through the air, and radiation to the ambient, as shown in Fig. 2(a). The heat transport is governed by Fourier’s law of conduction, as presented in Eq. (2). The steep tip–sample temperature gradient may challenge the applicability of Fourier’s law. However, when the mean free path () of the dominant heat carriers is much smaller than the tip radius (50 nm), and the SThM operates at slow scan rates (0.5–1 Hz, corresponding to ms per pixel), local thermal equilibrium is maintained.deshmukh2022direct; swoboda2023spatially. Therefore, Fourier’s law with an effective thermal conductivity can still be applied as a simple analytical model of heat transport. When the tip contacts the sample, a solid–solid tip–sample junction introduces an additional heat-conduction pathway, as described by Eq. (3). Based on the heat-transfer channels illustrated in Fig. 2(a,b), the relationships between the thermal resistances in non-contact and contact configurations are given by Eqs. (4) and (5), respectively.
The current applied to the probe through the integrated SThM power controller induces Joule heating in both the probe and the tip material. The probe current in CCM is higher than that in TCM. Under the constant-current condition of CCM, the generated electrical heat () flows from the tip into the sample. Therefore, the amount of heat transfer depends on the thermal conductivity of the sample. For highly thermally conductive materials, heat dissipates rapidly, resulting in a small temperature rise at the tip and a correspondingly low output voltage. In contrast, materials with high thermal resistance impede heat flow, causing heat to accumulate near the tip and producing a larger temperature rise and higher voltage output. The corresponding temperature change () can be expressed as shown in Eq. (6). Experimental verification of this relation is provided in Fig. S3 (Supplementary Material).
Using Eq. (6), the thermal signals can be converted into the corresponding tip temperature as follows:
Therefore, by substituting Eqs. (7) and (8) into Eqs. (2) and (3), we can obtain the relationship between the probe voltage and probe thermal resistance, as shown in Eq. (9):
Furthermore, since the tip radius (b) is very small, the associated thermal radius () is negligible relative to the sample size (>). Thus, under thermal equilibrium gu2002imaging, it is reasonable to approximate . Consequently, we obtain the following expression:
Therefore, using Eq. (10), the values of and can be expressed in terms of the measured thermal signals ( and ). Here, embodies the sample’s thermal transport characteristics, and its detailed formulation will be discussed in Sec. III.
III Methods
III.1 Sample preparation
A 5 mm × 10 mm Al film is patterned on a SiO2/Si substrate. A 6-inch CZ-prime wafer with a 1.5 m SiO2 film on its surface was diced into 12 mm × 12 mm pieces. The p-type (B doped) resistivity of the substrate is between 10-25 -cm. The surface was first cleaned by sequentially sonicating in toluene, acetone, and 2-propanol for 5 minutes, then rinsed with DI water. During the process, non-polar and polar contaminants were cleaned with toluene and acetone, respectively, and 2-propanol was used to remove acetone from the substrate. A positive photoresist (AZ1518) was spin-coated at 3000 rpm for 60 s with a dwell time of 5 s, which was then followed by a soft bake at 110 for 60 s. The substrate was exposed using a Heidelberg MLA150 maskless aligner, with the laser set to 405 nm and a dose of 160 mJcm-2. The sample was then developed in MF-26A for 35 s to remove the exposed photoresist over an area of 5 mm × 10 mm. A 15 nm Al film was evaporated by CHA e-beam evaporator. The evaporation was controlled by a Telemark 861 deposition controller, and the deposition rate was 2 Ås-1. A lift-off step was applied to remove the photoresist and Al from the unexposed area, leaving a 5 mm × 10 mm Al film on the surface.
III.2 Experimental setup and numerical data generation
The thermal signal data are obtained from the experiments conducted using an AFM (Dimension Icon® manufactured by Bruker Inc.) equipped with an integrated SThM module. The SThM probe is purchased from Bruker Corp. and it consists of a thin Pd line on Si3N4 cantilever, as shown in Fig. S5 (Supplementary Material). This probe is pre-mounted on the probe holder, and the probe tip serves as one of the arms of the Wheatstone bridge, as shown in the schematic in Fig. 1(a). Before each SThM scan, the Wheatstone bridge is balanced to minimize background thermal signal noise. Subsequently, an input voltage (> 0.5 V) is applied to the probe. The temperature increase at the SThM tip changes the resistance of the Pd heater, producing a voltage imbalance across the Wheatstone bridge. This imbalance is detected as the SThM voltage signal (). All measurements were conducted at a constant contact setpoint force and scan rate to ensure consistency. We used a gain of while extracting the change in voltage of the Pd heater. This gain helps amplify the corresponding thermal signals without any loss of information.
For thermal relaxation and steady-state measurements, the SThM tip rasters at a relatively low scan rate of 0.5–1 Hz (S2 supplementary material). This scan rate is maintained throughout the experiment to reduce variance in thermal signals due to thermal drift. The acquired images are analyzed using Gwyddion software, while the pixel-wise thermal signal data are extracted and processed through a specially developed Python code tailored for this study.
III.3 Measuring probe thermal resistance ()
Understanding the heat transfer between the probe and the surrounding system is essential for the quantitative measurement of thermal conductivity. The probe dissipates a finite amount of heat into the surrounding environment even in the non-contact state. Accurate quantification of this background signal is necessary to isolate the intrinsic thermal contribution from the sample. In earlier reports, the measurement of non-contact thermal resistance () was carried out by balancing the tip temperature using a thermal stage huang2024violation; deshmukh2022direct. This method, however, is prone to inaccuracies arising from parasitic thermal losses due to the roughness of the stage, non-uniform temperature profiles, and additional convective heat transfer to the probe.
We propose a circuit-based technique to measure the non-contact thermal resistance () by monitoring the probe-tip Joule heating power and the resulting temperature rise under an applied voltage. Unlike stage-based techniques, this approach eliminates dependence on thermal stage sensitivity and minimizes the influence of parasitic thermal effects from the environment. In this approach, Joule heating from the applied current raises the tip temperature. The temperature rise is directly proportional to the input power, where the proportionality constant corresponds to the non-contact thermal resistance, (Eq.12). The temperature rise () was calculated relative to the initial tip temperature, which was at room temperature ( K), and the corresponding initial electrical resistance of the tip ( 177 ). It is important to note that the probe electrical resistance (), which is output of balanced Wheatstone bridge, differs from the tip resistance () due to the presence of built-in limiting resistors in the probe, as shown in Fig. S4 (Supplementary Material). The applied power () was determined directly from Joule heating in the tip under the applied current using Eq.(11). The resulting data were fitted using linear regression, yielding an value of K/W as shown in Fig. 3. An instrumental uncertainty of 0.1 K in due to laser illumination was considered in the analysisspiece2019quantitative. It should be noted that the uncertainty associated with laser illumination may be relatively large; however, since the analysis focuses on the relative change in temperature (), the impact of this uncertainty is reduced. This result is consistent with previously reported values for similar probes, thereby validating the accuracy of the proposed measurement technique huang2024violation; razeghi2023single.
Here, denotes the voltage drop across the tip, extracted from the known SThM probe resistance and the Wheatstone bridge input voltage (Fig. 1(a)).
III.4 Modeling contact thermal resistance ()
As discussed in Sec. II, is an experimentally measured quantity that comprises the thermal properties of the sample. In addition to solid-solid thermal resistance (), also includes thermal resistances arising from radiation (), air conduction () and the water meniscus formed at the tip–sample contact (). The contribution of radiative heat transfer is often considered negligible, particularly under ambient SThM conditions with a cantilever–sample separation larger than 1 m li2025quantitative. The probe considered in our measurement has a tip height of 10 measured from the cantilever. In a humid environment, as shown in Fig. 2(b), a water meniscus forms at the tip–sample contact and serves as an additional pathway for heat transfer. A detailed study of water-meniscus formation and a theoretical model for heat transport through the meniscus were presented in Refs. gomes1999dc and luo1997sensor, respectively. Furthermore, Refs. zhang2020review and assy2015temperature reported that the contribution of the meniscus to the total tip–sample contact resistance () lies between 1 and 6. Therefore, the water-meniscus effect is treated as a 6 uncertainty in the values. Therefore, under the assumption that , we express the measured quantity as the sum of three thermal resistances, as shown in Eq. (13).
where:
: Tip-sample interface thermal resistance
: Spreading resistance (Yovanovich Eq. (14))yovanovich2002analytical; muzychka1999thermal
: Interface boundary resistance
: Geometric dilution factor
: Lateral spreading length
The impact of the interfacial thermal resistance () and its physically reasonable range are determined using FEM modeling, as discussed in Sec. IV. Furthermore, heat-spreading analysis is crucial for determining the thermal conductivity () of the sample. Yovanovich et al.yovanovich2002analytical; yovanovich2005four and Muzychka et al.muzychka1999thermal derived analytical solutions for the thermal spreading resistance (), as shown in Eq. (14), for a film on a substrate as a function of the film thickness () and the thermal conductivities of the film () and substrate ().
where is defined as , is the source radius, and is the first-order Bessel function of the first kind.
The buried interface thermal boundary conductance is a critical parameter in this analysis. However, as heat transfers through a thin film to the substrate, it spreads laterally before reaching the buried interface. The characteristic lateral spreading length () determines the effective area over which heat crosses the buried interface. For heat flow from a localized source of radius at the top of a thin film of thickness and thermal conductivity , deposited on a substrate with thermal conductivity , this spreading length is given by Eq. (15)muzychka2004thermal; yovanovich2003thermal; kennedy1960spreading.
This lateral spreading reduces the effective contribution of the buried interface resistance through a geometric dilution effect. The dilution factor represents the fraction of interface resistance “seen” by the heat flux and is given by Eq. (16).
Physically, heat spreading reduces the effective contribution of interface resistance by the ratio of the original contact area to the expanded area at the interface. As the spreading length increases relative to the contact radius, more heat bypasses the localized interface resistance, thereby reducing its impact on the total thermal resistance.
A critical insight is that both and depend on , which is the unknown quantity we seek to determine. This creates a self-consistency requirement: the spreading resistance and the effective buried interface resistance are both functions of the unknown film thermal conductivity. Consequently, the thermal resistance balance (Eq. 17)
must be solved iteratively to extract . This iterative solution converges to yield the thermal conductivity of the thin film, with the complete algorithm detailed in the section S2. (Supplementary Material).
IV Results and Discussion
IV.1 Finite-element multiphysics modeling of SThM probe
The variation in the working environment of SThM measurements and the uncertainty in the thermal resistance of the tip-sample interface () make quantitative measurements challenging. Therefore, a fundamental understanding of thermal transport at the tip-sample interface is essential for the quantitative determination of local thermal conductivities of the sample. In this study, a detailed three-dimensional finite element model (3D-FEM) is developed to evaluate the interfacial thermal resistance () by simulating the heat transfer mechanisms occurring at the probe–sample junction. Accurate determination of provides insights into the complex heat transfer pathways occurring between the tip and the sample. This 3D-FEM model is developed using COMSOL Multiphysics® to systematically explore the effects of multiple physical parameters on the tip–sample thermal interaction. Fig. 4(a) illustrates the SThM probe and sample structure, along with their meshed geometry. The probe comprises a Si3N4 cantilever, an Au pad connected to the tip, and a Pd tip. Detailed dimensions and structural specifications of probe are provided in Fig. S5 (Supplementary Material). The model incorporates two primary physical processes: (1) heat transfer in solids and (2) electromagnetic heating (Joule heating in the Pd resistor). All exposed surfaces are set as adiabatic, except for the interface between the tip and the sample.
The quasi-equilibrium heat transfer equation with internal Joule heating is employed to model heat transport throughout the system:
where is the thermal conductivity. The volumetric Joule heat source arises from electrical biasing of the Pd thermistor and is given by
Where is the applied voltage, is the temperature-dependent electrical resistance of the Pd thermistor, and is the effective heater volume. This expression corresponds to uniform distribution of electrical power dissipation within the resistive element. Convection term was not considered in Eqs. (18) due to its negligible contribution under the given experimental conditions zhang2023realizing; bodzenta2022scanning; deshmukh2022direct.
The heat generated at the probe tip is transferred through the tip–sample interface and subsequently into the sample. To accurately capture the variation in , the SThM tip–sample contact is modeled over a wide range of values, from to K/W, as illustrated in Fig. 5(a). This simulation range is selected based on experimentally obtained thermal resistance values for nanoscale contactsli2025quantitative; gomes2015scanning. At low interfacial resistance ( K/W), the contribution of to the overall thermal resistance () is negligible, indicating perfect heat transfer across the interface. On the other hand, at high interfacial resistance ( K/W), heat transfer between the tip and sample becomes significantly suppressed, effectively isolating the thermal transport pathways. The resulting probe temperature profiles for different values are presented in Fig. 5(a). At high interface thermal resistance values (), the sample temperature remains constant at room temperature, indicating the absence of local heating. Therefore, cannot exceed , since local heating of the sample is need for thermal resistance measurement.
The effect of the on local sample heating was examined for two materials with strongly contrasting intrinsic thermal conductivities: Cu (300 W/mK) and SiO2 (1.38 W/mK). As shown in Fig. 4(b–d) for SiO2 and Fig. 4(f–h) for Cu, identical values of result in qualitatively similar surface temperature distributions beneath the tip, despite the large difference in bulk thermal conductivity. Because the COMSOL model considers smooth surfaces for all domains, the resulting temperature profiles for Cu and SiO2 are expected to be identical, which is in agreement with previous studies showing that is independent of the underlying material in the absence of surface roughness effects li2025quantitative; pernot2021frequency; bahrami2004thermal. At lower interfacial thermal resistance ( K/W), efficient heat transfer from the probe leads to a broader thermal diffusion profile in both materials. In contrast, increasing K/W strongly suppresses heat flow across the interface, resulting in negligible local surface heating for both Cu and SiO2. Importantly, the experimentally measured total thermal resistance lies in the range K/W and is accompanied by clear local surface heating beneath the probe. Therefore, must be K/W and cannot exceed the experimentally inferred thermal resistance.
Furthermore, the difference between the SThM tip temperature and the corresponding local sample temperature variation is plotted in Fig. 5(b). The change in tip and sample temperature () is shown for various values. The results shown in Fig. 5(a,b) demonstrate that plays a dominant role in the quantitative determination of the sample’s local thermal conductivity. While is treated parametrically in the FEM simulations to examine its influence on the measured temperature response, this approach is adopted to establish a physically constrained relationship between the experimentally accessible temperature contrast and . By comparing the experimentally measured (Fig. 3) with the trends predicted by the FEM model shown in Fig. 5 (b), a plausible and bounded range of is identified for the given probe–sample contact conditions.
IV.2 Experimental determination of model parameters
To experimentally validate the quantitative SThM model, a 15 nm aluminum (Al) thin film deposited on a SiO2 substrate was selected as the test sample. The film was patterned adjacent to the bare SiO2 substrate, enabling direct comparison between regions of differing thermal conductivity under identical measurement conditions. The sample preparation and fabrication details are provided in the Sec. III (A). All SThM measurements were carried out at a controlled room temperature of about 294 K, monitored continuously by a temperature regulation system. The SThM probe was positioned near the Al/SiO2 interface and brought into mechanical contact with the sample surface. This contact led to a sudden reduction in probe temperature, indicating the activation of an additional thermal conduction pathway at the tip–sample junction, thereby enhancing heat dissipation into the sample (see Fig. S1 from the Supplementary Material). Subsequently, as the tip scanned the sample area, the probe recorded thermal signals from the sample as temperature variations, which allowed the calculation of the sample’s thermal resistance. A total of 7 independent SThM scans were acquired to ensure the reproducibility of the measurements. The corresponding pixel-level thermal signal and topography data were extracted using the SPM image analysis software named Gwyddion. The data were processed and analyzed using a custom-developed Python script designed specifically for quantitative pixel-wise thermal signal correlation and noise reduction. This in-house code enabled automated data extraction, normalization, and statistical comparison across multiple scans, significantly improving the accuracy and consistency of the thermal mapping results.
Figure. 6(a,c) shows the topography and thermal signal map of 15 nm-thick Al thin film patterned on a SiO2 substrate. A strong thermal signal contrast was observed between the Al film and the SiO2 substrate, reflecting the significant differences in their thermal conductivities. Two approach-retraction curves for either side of the interface are used for the conversion between the SThM signal and thermal resistance using Eq. (10). Further details about the approach-retraction curve are provided in Fig. S1 (Supplementary Material). The relative change in the thermal signal (20 mV) from the SiO2 surface to the Al film corresponds to a decrease in thermal resistance from the SiO2 surface to the 15 nm-thick Al film. Therefore, this decrease in thermal resistance indicates a higher thermal conductivity of the Al film compared to the SiO2 substrate.
To extract the thermal resistance associated with Al and SiO2, a pixel-to-pixel correlation between the height and thermal signal maps is performed. This correlation is used to construct a 2D histogram (see Fig. S2 from the Supplementary Material), which enables a separation of the thermal responses from Al and SiO2 surfaces. Traditional analysis methods often rely on arbitrarily selected line-cuts or cropped regions of interest, introducing potential user bias in the selection process. In contrast, the present approach systematically processes all available data points from the thermal signals measurement, ensuring reproducibility and minimizing subjective interpretation. This comprehensive analysis leverages the full spatial information content, thereby improving statistical confidence in the extracted thermal properties. Subsequently, the acquired data are subjected to statistical processing to extract thermal properties. However, this analysis is susceptible to errors arising from experimental artifacts—such as surface contamination, tip-sample contact variations, or instrumental noise—which can distort the thermal signal and compromise measurement accuracy. Therefore, to quantitatively extract representative thermal resistance values while accounting for spatial heterogeneity, a 2D histogram of thermal resistance on the y-axis and film thickness on the x-axis is constructed (Fig. S2 of Supplementary Material). This histogram is then projected onto the thermal resistance axis to obtain statistically representative values. This projection combines measurements across all thickness values, yielding a one-dimensional distribution from which statistically robust resistance values can be determined. The resulting 1D distribution is modeled using a double-Gaussian function, from which the mean peak positions corresponding to the thermal resistances of Al and SiO2 are extracted. The statistical uncertainties in these mean values are obtained from the covariance matrix of the nonlinear least-squares fitting procedure. From this analysis, the mean thermal resistances were found to be 2.60 0.03 MK/W and 5.41 0.04 MK/W for Al and SiO2, respectively.
The measured thermal resistances MK/W and MK/W provide the experimental foundation for quantitative determination of the thermal conductivity of the Al thin film. Following the theoretical framework established in Sec. III, we model the measured contact-to-noncontact thermal resistance difference as a series combination of three contributions: the tip-sample interfacial thermal resistance (), the sample spreading resistance (), and the Al/SiO2 buried interface thermal boundary resistance (), as given by Eq. (13). This formulation enables direct extraction of the Al thermal conductivity from the experimental data through an iterative algorithm (Eq. 17) that accounts for the interdependence of and on the unknown film thermal conductivity.
IV.2.1 Tip Interface Resistance Extraction
As shown in Fig. 6(e), the experimentally measured values are K/W. Therefore, it can be concluded that is also K/W. In addition, the tip–sample contact resistance () is mainly governed by surface roughness, the applied tip–sample normal force, and the resulting contact area pernot2021frequency; bahrami2004thermal; puyoo2010thermal. Therefore, the thermal contact between the SThM probe and the sample may slightly differ for Al and SiO2 surfaces; however, it is very challenging to determine the exact for each case. Therefore, for simplification, we assumed the same for both Al and SiO2 interfaces. This assumption introduces negligible uncertainty in the measurement because both surfaces exhibit comparable roughness (see Fig. S6 in the Supplementary Material). Additionally, the tip-sample normal contact force was maintained with an AFM setpoint value of 0.6 V, which corresponds to < 10 nN contact force for this type of probes. Therefore, the tip-sample interface resistance is extracted from a reference measurement on bare SiO2/Si (no thin film, ).
where is calculated using Eq. (14) with known (1.38 W/mK) and (150 W/mK). The parameters used in this analysis are provided in Table S1 (supplementary material).
IV.2.2 Iterative Algorithm for Extraction
Literature values for metal/dielectric interfacesstoner1993kapitza; lombard2014influence; schmidt2008pulse; cheaito2015thermal span a considerable range. For the Al/SiO2 interface, we adopt a central value of MW/(m2K) with a realistic range of –250 MW/(m2K) to account for this uncertainty hopkins2010criteria; kwon2021thermal; zhu2010ultrafast.
Applying the iterative algorithm to the experimental measurements yields a thermal conductivity of W/(mK) at the central values of MW/(m2K) and nm. See Section S3 (supplementary material) for the detailed calculation. In our 15 nm Al film, a sub-stoichiometric Al2O3-x layer (1–2 nm thick) forms naturally on the Al surface due to exposure to ambient airyang2018liquid; gorobez2021growth, reducing the effective conductive metallic thickness to approximately 13–14 nm. A comprehensive uncertainty analysis accounting for variations in both interface conductance (–250 MW/(m2K)) and film thickness (–15 nm) produces asymmetric error bounds of W/(mK), yielding a final result of W/(mK). This represents a 5.3-fold reduction from bulk aluminum (237 W/(mK)), consistent with Fuchs-Sondheimer theoryfuchs1938conductivity; Sondheimer1952MeanFreePath for films of thickness comparable to the electron mean free path (). The thermal resistance budget (Table 2) reveals that the tip-sample interface dominates the total measured resistance (70.8%), with spreading resistance contributing 26.2% and the diluted buried interface resistance only 3.1%. The geometric dilution factor of indicates that lateral heat spreading over a characteristic length of nm substantially reduces the effective contribution of the Al/SiO2 interface resistance. A complete matrix of extracted thermal conductivity values for all combinations of and thickness is provided in the Table. S1 of Supplementary Material.
| Component | Value (MK/W) | Percentage | Description |
| (tip-Al) | 1.84 | 70.8% | Tip-sample interface |
| (Al/SiO2) | 0.68 | 26.2% | Spreading resistance |
| (Al/SiO2) | 0.08 | 3.1% | Diluted |
| Total | 2.60 | 100% | = |
| Metal | (W/mK) | (W/mK) | (W/mK) |
| Al | 6 | 246 | 237–252 |
Data adapted from Refs. jain2016thermal; kittel2018introduction.
Discussion
The thermal conductivity of Al consists of both phonon () and electron () contributions, with electrons typically dominating heat transport in bulk. Jain et al. performed detailed first-principles calculations to examine both phonon-phonon and electron-phonon scattering processes in bulk Al, quantitatively determining the individual contributions of and to the overall thermal conductivityjain2016thermal. Furthermore, they quantified the mean free path () distributions of both phonon and electron heat carriers by analyzing their respective thermal conductivity accumulation functions at 300 K (see Fig. S7 of Supplementary Material). Analysis of the thermal conductivity accumulation function demonstrates that the electron contribution decreases substantially as decreases, indicating that electrons with short contribute less to heat transport than those with longer .
The accumulation function data at 300 K indicate that phonon thermal conductivity in Al is dominated by carriers with mean free paths ranging from 1 nm to 22 nm (comprising 90% of ), while electron thermal conductivity requires longer mean free paths of 10 nm to 25 nm for similar accumulation (90% of ). The onset of contributions also differs markedly: phonon transport begins at nm, whereas electron transport only becomes substantial at nm. These characteristic length scales directly impact thermal transport in confined geometries. This confinement preferentially suppresses electron transport, as the film dimension is smaller than the typical electron mean free paths, resulting in diffuse boundary scattering that significantly reduces .
Fig. 7 compares the SThM-extracted thermal conductivity of the 14 nm Al film ( W/(mK)) with literature values for polycrystalline Al thin films of varying thickness measured by different techniques, including frequency-domain thermoreflectance (FDTR) Schmidt2010ThinMetalFDTR, photothermal mirage Wu1993PhotothermalThinFilms, electrical heating with temperature profile analysis lugo2016thermal, and electron diffraction thermometry Volkov1976AluminumThermal. Boltzmann transport equation (BTE) modeling results from Jin et al. Jin2008ElectronMeanFreePath are also included. Our result falls on the expected trend of decreasing thermal conductivity with decreasing film thickness, consistent with enhanced electron scattering at surfaces and grain boundaries.
The dashed curve in Fig. 7 is a trend line that captures the thickness-dependent behavior and remains consistent with the combined Fuchs-Sondheimer and Mayadas-Shatzkes electron scattering model fuchs1938conductivity; Sondheimer1952MeanFreePath; mayadas1970electrical. Using the literature value of nm for the electron mean free path in aluminum Gall2016, this theoretical framework predicts W/(mK) for a 14 nm film (Supplementary Section S4), which falls within our experimental uncertainty range. The spread in literature data reflects variations in film microstructure, deposition conditions, and measurement techniques across different studies. The agreement between our SThM measurement, the theoretical prediction, and existing literature data validates the iterative extraction methodology developed in this work.
The thermal signal and topography appear elevated along the Al/SiO2 boundary (see Fig. 6(a, c)). The elevated topography introduces a minor artifact in the thermal signal due to AFM feedback sensitivity effects at the step height. Consequently, the higher thermal resistance extracted at the edge arises from the same artifact, as it is directly derived from the thermal signal. Nevertheless, Fig. 6(e) demonstrates the significant potential of this approach for probing thermal transport variations across dislocation defects and grain boundaries with sub-100 nm spatial resolution. Furthermore, validation using established approaches, such as those described in Refs. tiwari2013anisotropic; isotta2023microscale for assessing thermal property variations across grain boundaries with different densities, would be highly beneficial, particularly given the high spatial resolution offered by SThM.
V Conclusion
We experimentally quantified the thermal resistance of the SThM probe using a circuit-based calibration method and employed FEM modeling to investigate the underlying heat transfer mechanisms. The contributions of various probe–sample heat transfer channels, along with their corresponding measurement strategies, were systematically examined. Furthermore, the thermal conductivity of a 15 nm Al thin film, deposited by e-beam evaporation on a few-micrometer-thick SiO2 substrate, is measured. The robustness of the pixel-to-pixel correlation approach for extracting representative thermal signals is discussed. In addition, statistical and measurement uncertainties in thermal resistance are quantified, considering the underlying assumptions. The measured thermal conductivity of the Al thin film is in good agreement with the theoretical frameworks and the published experimental data.
Overall, as demonstrated in this study, the SThM-based thermal metrology approach enables quantitative and spatially resolved characterization of thermal transport at the nanoscale. This technique provides sub-100 nm lateral resolution, allowing direct mapping of local heat conduction pathways with precision far beyond that achievable by conventional optical methods such as TDTR and FDTR. Unlike TDTR and FDTR, which are fundamentally constrained by optical diffraction limits and thermal penetration depth—thereby limiting their reliability for ultrathin films—SThM imposes no such thickness restrictions. This advantage makes it particularly effective for studying two-dimensional materials, nanoscale heterostructures, and complex interfaces. SThM emerges as a powerful and complementary tool for nanoscale thermal transport characterization, offering insights into defect-, boundary-, and interface-driven effects that are inaccessible through traditional macroscopic or optical approaches.
Supplementary Material
See the supplementary material for the raw data that supports the main findings of this study.
Acknowledgements.
We acknowledge Dr. Amy Marconnet, Professor of Mechanical Engineering at Purdue University, for her valuable guidance and for providing access to the COMSOL license.Data Availability Statement
The experimental and modeling data that support the findings of this study are available from the corresponding author upon reasonable request.